Calculate the number of mols of copper II sulphate in 50.0cm3 of 1.00moldm3 solution
1
Expert's answer
2020-04-18T07:08:42-0400

According to the equation:

M = n / V

where M - molar concentration, n - number of moles, V - volume.

From here:

n = M × V

As 1.00 mol/dm3 = 1.00 mol/L and 50.0 cm3 = 50.0 mL, the number of moles of copper II sulphate equals:

n(CuSO4) = M(CuSO4) × V(CuSO4) = 1.00 mol/L × 50.0 mL = 1.00 mol/L × 50.0 × 10-3 L = 0.0500 mol


Answer: 0.0500 mol